1. An object is 70 micrometers long and 47.66 micrometers wide.how long and wide is the object in km?
1
Expert's answer
2021-06-03T18:26:50-0400

(a)

"l=70\ \mu m\cdot\dfrac{10^{-6}\ m}{1\ \mu m}\cdot\dfrac{1\ km}{10^3\ m}=7.0\cdot10^{-8}\ km."

(b)

"w=47.66\ \mu m\cdot\dfrac{10^{-6}\ m}{1\ \mu m}\cdot\dfrac{1\ km}{10^3\ m}=4.7\cdot10^{-8}\ km."